Week of 170911

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Kate (DB, chair), Ulf (NDGF), Vincent (security), Vladimir (LHCb), Eddie (monitoring), Maarten (Alice, WLCG), Andrea M (storage, MW), Gavin (comp), Marian (network)
  • remote: Andrzej O (Atlas), Onno (NL-T1), Chih-Sun (ASGC), David B (IN2P3), Di Qing (TRIUMF), John K (RAL), Kyle (OSG), Luca L (CNAF), Sang Un (KISTI), Xin Zhao (BNL), Christoph W (CMS), Dave M (FNAL),

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Normal activities, currently running 350k jobs with peaks up to 700k jobs using NERSC_Cori
      • Increased limit of data overlay jobs to 1600, need to watch Frontier services
    • Problems
      • Network issues at CERN affected CERN CEPH instance, EOS, monitoring and other dependent services
      • RAL Frontier overloaded on weekend by group production with lots of data taken from DB. Production stopped. But service needed 2 reboots on Sta/Sun to come back to operation.
Maarten asked if it is possible that frontier servers are not powerful enough in RAL. John commented they are 5 machines behind loadbalancer. Andrzej remarked that is issue is being followed up by a Frontier expert.

  • CMS reports ( raw view) -
    • Continued high activity ~160k CPU core utilized in Global Pool during last days
    • Still high pressure on transfer system
    • Recent T1 related operational issues
      • Castor based SE at RAL seems full GGUS:130467, might need adjustment of DDM thresholds
      • Access failures at a known I/O intense sample at PIC GGUS:130441
      • CMS EOS instance dropped out of the CMS data federations GGUS:130459
      • Bad transfer quality from CERN Castor GGUS:130363
      • Some time outs for transfers from FNAL GGUS:130380
Kate will follow up on the federation issue. Dave will update the FNAL issue.

  • ALICE -
    • Continued very high activity
    • CERN: EOS-ALICE not reachable from outside Fri morning (OTG:0039654)
    • CERN: Oracle cloud project resources being used since the weekend
Huge thanks from openlab Oracle to both Maarten and Ben Jones for the work done on the cloud.

  • LHCb reports ( raw view)
    • Activity
      • Monte Carlo simulation, data processing and user analysis
    • Site Issues
      • T1:
        • Failed transfers from IC to SARA (IPV6) (GGUS:129946); "Geant have confirmed that they are unable to ping mouse1.grid.sara.nl from geant-lhcone-gw.mx1.lon.uk.geant.net"
        • Access file problem at GRIDKA (GGUS:130478)

Sites / Services round table:

  • ASGC: ntr
  • BNL: dCache srm became unresponsive for a couple of hours last Thursday, due to intensive srm DB cleanup, recovered shortly.
  • CNAF: ntr
  • EGI: nc
  • FNAL: EOS was upgraded to the version in-sync with CERN.
  • IN2P3: Scheduled maintenance on next Tuesday September 19th: CEs will be down and back during the afternoon, SEs will be down in the morning. Downtimes will be declared by end of the day.
  • JINR: Some problems in routing with LHCONE because of an error when reconfiguring routing tables. Fixed.
  • KISTI: ntr
  • KIT: nc
  • NDGF: On Monday last week after the meeting a dcache pool with 70T of alice disk data developed problems. After some moving around to make place for a rescue mission, we got it mounted RO on Friday. So far 15T of data has been saved, the process is ongoing.
  • NL-T1:
    • Last Thursday around 0:00 CEST, a network outage brought down the SARA-MATRIX SRM. It was fixed Thursday morning around 10:30. It was an end-of-row switch that had crashed; the vendor is investigating why that happened.
  • NRC-KI: nc
  • OSG: ntr
  • PIC: nc
  • RAL: CMS filled their Castor disk over the weekend. Also there were problems with the Atlas frontier machines.
  • TRIUMF: NTR

  • CERN computing services: ntr
  • CERN storage services:
    • FTS: On Wed 7th at ~14:00 we noticed a huge number of links submitted to FTS prod, ( ~2.5.k compared to ~1.5k which we normally have) which caused the Transfer Optimizer to work very slowly, so basically the links have been not optimized during Wed 7th on the afternoon. This has been reported to the experiments and ATLAS confirmed that due to a bug their side an unplanned number of submissions has been sent to our FTS. The problem got fixed during the evening. Based on this incident we are planning to investigate how to improve the performances of the Optimizer to cope with higher number of links submitted.
  • CERN databases:
    • OS updates application to DB hosts (rolling intervention)
      • CMSONR: Monday, September 18th at 10:00 - 12:00
      • CMSONR ADG and CMSARC: Tuesday, September 19th at 9:30 - 10:30
      • CMSR: Tuesday, September 18th at 10.30 - 12.00
      • ALICESTG/ATLASSTG/LHCBSTG Monday, September 18th at 9:00 - 12:00
      • CMSSTG Wednesday, September 20th at 9:00 - 12:00

  • GGUS: ntr
  • Monitoring: NTR
  • MW Officer: NTR
  • Networks: LHCONE to be involved in Sara issue.
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2017-09-11 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback